# Spherical orbits around Kerr-Newman and Ghosh black holes

### v4

@@ III.1.4 Selected Orbits
Using the analytical solutions $(\theta,\phi,t)$ we will show typical of orbits for non-extreme, extreme, and naked conditions of regular rotating black holes. Unlike the KN case, however, here we have an additional parameter to consider, $k$, which appears non-linearly in the Eq. (58),
@@
Non-extremal
  Orbit   $Mk$   $r_{SO}/M$   $C/M^{2}$   $u_{0}$   $\Delta\phi$
  (a)   0   2.892   24.412   0.088   22.436
  (b)   0.1   2.743   22.524   0.088   23.698
  (c)   0.5   1.996   14.245   0.087   33.482
Table 7: SPO Parameters around a non-extremal Ghosh black hole with $\Phi=M$ and $a/M=0.3$.
The non-extremal condition is not compatible with $Mk>>1$ regime. There exists no parameter $a/M$ that satisfies:
  1. 1.
    $Mk>>1$,
  2. 2.
    $r_{SO}>r_{+}\neq r_{-}$, and
  3. 3.
    Eq (70),
simultaneously. For example at $a/M\sim 10^{-6}$ the maximum $Mk$ is $\sim 0.73$. In this work, we choose $a/M=0.3$, and the maximum $Mk$ we can have here is $Mk\sim 0.5$. In Fig. 11, we depict three examples of orbits with $\Phi=M$ in the $(x-y-z)$, the $x-y$, and $x-z$ planes, respectively. The radius and other parameter values of the orbit are provided in Table7. The SPOs appear predominantly equatorial, characterized by minimal latitudinal oscillations with elongated oscillation periods. This observation is consistent with expectations for the slowly rotating black hole regime ($10<a/M<$ 1).
Figure 11: Spherical photon orbits around non-extremal Ghosh black hole with $\Phi=M$ in the $(x-y-z)$ as well as the projective $x-y$ and $x-z$ planes, respectively. The direction of the black hole’s rotation is from west to east. Figures (a), (b), and (c) are cases with $Mk=0$, $Mk=0.1$, and $Mk=0.5$, respectively.
@@
In the extremal Ghosh black hole case, we also attempted to identify the photon boomerang, characterized by the condition where Eq. 75 equals $\pi$. Through our numerical investigation, we found only an approximate photon boomerang with $k=0.05$. We label it as “approximate” because the value of $\Phi=0.07$ is non-zero. This is detailed in Table 10, and the corresponding trajectory is depicted in Fig. 14.
@@
Certain combinations of ${a/M,\ Mk}$ in Eq.(77) can give rise to complex solutions for $r_{\pm}$. In the context of a conventional black hole, such instances correspond to what is termed a ”naked singularity,” a phenomenon deemed unstable and unphysical according to the cosmic censorship theorem Penrose:1964wq . However, within the framework of a regular black hole, the concept of singularity vanishes. This scenario is often referred to as ”the no-horizon spacetime” in the literature. While discussing trajectories around a naked singularity is nonsensical, the notion is perfectly acceptable in the case of a regular no-horizon spacetime.
@@
We illustrate orbit trajectories around the no-horizon solution in Figs. 15-16, delineating parameters detailed in Table 11. Each figure represents distinct values of $\Phi$ and $k$, as the stability requirement imposes stringent constraints on them. Given the negative $\Phi$ values, the occurrence of the Lens-Thirring effect is inevitable. This phenomenon exerts a pronounced influence on orbits near the equator compared to those near the poles. Notably, orbits with negative $\Phi$ values maintain a prograde motion around the equator, transitioning to a retrograde direction as they approach the poles. This observation holds particular significance for orbits with $\Phi\sim 0$, as the Lens-Thirring effect possesses the capability to counteract the photon’s angular momentum.
@@ IV Conclusion
For stable spherical timelike orbits (STOs), both KN and Ghosh BHs allow orbits to exist outside their horizons. In the KN scenario, bound orbits occur when $E^{2}<1$, while $E^{2}=1$ signifies a marginally bound orbit. Unlike null orbits, timelike orbits do not exhibit retrograde-prograde switching. In the Ghosh case, higher values of $k$ drive trajectories closer to the pole, resulting in larger latitudinal oscillation amplitudes. Notably, the concept of innermost stable spherical orbits (ISSOs) in spherical trajectory parallels the innermost stable circular orbits (ISCOs) in equitorial motion. The ISSO radius in KN can be determined by solving Eqs. (53)-(54). These solutions, depicted in Fig. 8, manifest as two branches of $r_{ISSO}\left(C\right)$. A typical ISSO trajectory in the KN BH, illustrated in Fig. 9, appears densely concentrated around the equator initially, gradually filling the entire solid angle as it completes one latitudinal oscillation. In contrast, ISSO behavior in the Ghosh BH presents intriguing nuances. Unlike the KN case, the nonsingular nature of the Ghosh metric causes ISSO branches to terminate at finite values, as depicted in Figs. 20-22. A representative ISSO trajectory is shown in Fig. 23. As a closing statement in this paragraph we should briefly comment on this ISSO issue. It is evident that in both types of black holes under examination, the radial potential $R(r)$ is at least quartic in $r$. This leads to two roots for each solution of $R^{\prime\prime}=0$, hence the presence of two branches. The lower branch in each $r_{ISSO}$ vs $C$ plot should be the one that represents the ”true” ISSO, as the name by itself stands.

### v6

@@ III.1.4 Selected Orbits
Using the analytical solutions $(\theta,\phi,t)$ we will show typical orbits for extreme, and naked conditions of regular rotating black holes. Unlike the KN case, however, here we have an additional parameter to consider, $k$, which appears non-linearly in the Eq. (59),
@@
The extremal condition is not compatible with $Mk>>1$, similar to the non-extremal case. There exists no parameter $a/M$ that satisfies:
  1. 1.
    $Mk>>1$,
  2. 2.
    $r_{SO}>r_{+}\neq r_{-}$, and
  3. 3.
    Eq (71),
@@
In the extremal Ghosh black hole case, we also attempted to identify the photon boomerang, characterized by the condition where Eq. (76) equals $\pi$. Through our numerical investigation, we found only an approximate photon boomerang with $k=0.05$. We label it as “approximate” because the value of $\Phi=0.07$ is non-zero. This is detailed in Table 7, and the corresponding trajectory is depicted in Fig. 11.
@@
Certain combinations of ${a/M,\ Mk}$ in Eq.(78) can give rise to complex solutions for $r_{\pm}$, which, in a conventional black hole, is termed naked singularity. However, within the framework of a regular black hole, the concept of singularity vanishes. This scenario is often referred to as ”the no-horizon spacetime” in the literature. While discussing trajectories around a naked singularity is nonsensical, the notion is perfectly acceptable in the case of a regular no-horizon spacetime.
@@
We illustrate orbit trajectories around the no-horizon solution in Figs. 12-13, delineating parameters detailed in Table 8. Each figure represents distinct values of $\Phi$ and $k$, as the stability requirement imposes stringent constraints on them. Given the negative $\Phi$ values, the occurrence of the Lens-Thirring effect is inevitable. This phenomenon exerts a pronounced influence on orbits near the equator compared to those near the poles. Notably, orbits with negative $\Phi$ values maintain a prograde motion around the equator, transitioning to a retrograde direction as they approach the poles. This observation holds particular significance for orbits with $\Phi\sim 0$, as the Lens-Thirring effect possesses the capability to counteract the photon’s angular momentum.
@@ IV Conclusion
For stable spherical timelike orbits (STOs), both KN and Ghosh BHs allow orbits to exist outside their horizons. In the KN scenario, bound orbits occur when $E^{2}<1$, while $E^{2}=1$ signifies a marginally bound orbit. Unlike null orbits, timelike orbits do not exhibit retrograde-prograde switching. In the Ghosh case, higher values of $k$ drive trajectories closer to the pole, resulting in larger latitudinal oscillation amplitudes. Notably, the concept of innermost stable spherical orbits (ISSOs) in spherical trajectory parallels the innermost stable circular orbits (ISCOs) in equitorial motion. The ISSO radius in KN can be determined by solving Eqs. (54)-(55). These solutions, depicted in Fig. 6, manifest as two branches of $r_{ISSO}\left(C\right)$. A typical ISSO trajectory in the KN BH, illustrated in Fig. 7, appears densely concentrated around the equator initially, gradually filling the entire solid angle as it completes one latitudinal oscillation. In contrast, ISSO behavior in the Ghosh BH presents intriguing nuances. Unlike the KN case, the nonsingular nature of the Ghosh metric causes ISSO branches to terminate at finite values, as depicted in Figs. 17-19. A representative ISSO trajectory is shown in Fig. 20. As a closing statement in this paragraph we should briefly comment on this ISSO issue. It is evident that in both types of black holes under examination, the radial potential $R(r)$ is at least quartic in $r$. This leads to two roots for each solution of $R^{\prime\prime}=0$, hence the presence of two branches. The lower branch in each $r_{ISSO}$ vs $C$ plot should be the one that represents the ”true” ISSO, as the name by itself stands.